OpenStudy (anonymous):

Evaluate the expression 1024^1/5

OpenStudy (anonymous):

\[1024^1/5\]

OpenStudy (jhannybean):

\[\large 1024^{1/5}\] is this it?

OpenStudy (anonymous):

yes that is what i need to evaluate. if you could also please explain how you do it i will be thankful and give you a medal :)

OpenStudy (jhannybean):

Sure thing.

OpenStudy (anonymous):

thanks :)

OpenStudy (jhannybean):

\[\large \sqrt[5]{1024}\]\[\large \sqrt[5]{256 \cdot 4}\]\[\large \sqrt[5]{64 \cdot 4\cdot 4}\]\[\large \sqrt[5]{4 \cdot 4 \cdot 4 \cdot 4 \cdot 4}\] So for every 5 numbers of the same type, only 1 will come out of the square root. Therefore, what will our answer be?

OpenStudy (anonymous):

i don't get it. so im not allowed to take all the 4's out from the square root into a 2 only 1 on them?

OpenStudy (jhannybean):

the number outsideof the square root, "5" tellsyou how many pairs ofthe exact number, inthis case 4, are taken out of the square root. It's like asking a child to take out 5 pairs of red colored balls from a jar with an infinite amount of different colored balls.

OpenStudy (anonymous):

so would the answer be 10, like as if you take out 5 4's and make them each 2 which equals 10?

OpenStudy (jhannybean):

\[\huge \sqrt[\color{red}5]{4 \cdot 4\cdot 4 \cdot 4 \cdot 4}\] that highlighted portion right there. It's basically asking you for 5 of the same number. The answer would be 4 since for every 5 fours, we take 1 four out from the square root.

OpenStudy (anonymous):

OHHHHHH. srry it took so long for me to get it. you were very helpful and patient, thank you! :)

OpenStudy (jhannybean):

no problem! :)